Write the multiplicative inverse of -13/19 and -7
Write the multiplicative inverse of -13/19 and -7
Write the multiplicative inverse of -13/19 and -7

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear student

-13/19 = -19/13

-7 = -1/7


Thanks
Yashraj
3 years ago
  1. Practice previous year questions many times..
  2. Appear for a test series and analyse it properly (Giving 2–3 hours)
  3. Revise your notes multiple times till they enter your soul :)
  4. Find your strong points and chapters, work on them..
  5. Solve other people’s doubts. It makes your own concepts very clear